OpenStudy (anonymous):

What is the solution of log2x − 4 16 = 4? x = −3 x = 3 x = −2 x = 2

OpenStudy (anonymous):

\[\log_{2x-4} 16=4\] I keep getting x=1/2

OpenStudy (anonymous):

first get the log by itself on the left side

OpenStudy (anonymous):

do i just go about solving 2x-4? 2x+4 - 4+4 2x-0 2/2? I'm not sure lol

OpenStudy (anonymous):

no no solve just like you would algebraically. so add 416 to both sides

OpenStudy (anonymous):

No, I wrote it underneath my post. It's log(2x-4) 16=4

OpenStudy (anonymous):

It isn't 416............................

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

Thanks.. are you sure?

OpenStudy (anonymous):

yes sir.

OpenStudy (anonymous):

I'm a girl oo"

OpenStudy (anonymous):

yes ma'am